# TrustLLM: Smelling and Reasoning Smart Contract Vulnerability

### v1

@@ IV-A Experimental Setup
In our study, we carefully selected a series of benchmark models, categorized into two groups: LLMs for zero-shot learning and pre-trained code models based on fine-tuning, to ensure a comprehensive and sound comparative analysis. For zero-shot learning LLMs, we chose CodeLlama-13b-Instruct, CodeLlama-34b-Instruct [49], GPT-3.5, and GPT-4 as benchmarks, representing the current state-of-the-art. Additionally, we selected CodeBERT [23], GraphCodeBERT [56], CodeT5 [28], UnixCoder [57], and CodeLlama-13b [49] to train classifiers. Among these, CodeBERT, GraphCodeBERT, CodeT5, and UnixCoder underwent a complete model fine-tuning process to adapt to the specific code classification task. In particular, CodeLlama-13b employs LoRA for lightweight tuning and uses the last token representation for classification. Note that that our method is different; TrustLLM’s Detector achieves classification by generating label names as task outputs.
@@ IV-B Research Questions (RQs)
How does the performance of TrustLLM in detecting vulnerabilities compare to other models? This question aims to understand how the effectiveness of Detector in detecting vulnerabilities compares to that of other existing models. The focus is on comparative analysis, involving metrics accuracy, precision, recall, and F1 score, to evaluate and contrast the performance.
@@
To what extent do the explanations generated by TrustLLM’s Reasoner align with the real reasons? RQ2 concerns the quality of the explanations the Reasoner provided for the decision of the Detector. It questions whether the reasons given by TrustLLM correspond to the actual reasons behind the vulnerabilities, emphasizing the interpretability and trustworthiness of the model.
@@
How does TrustLLM compare with an integration model that performs detection and reasoning simultaneously? Our method is based on a generative model, with two models trained on the generated labels and reasons, respectively. Another approach uses a single model to generate both reasons and labels. This question explores the effectiveness and impact of integrating the Detector and Reasoner components into one.
@@
Besides the RQs above, we also used our model to audit two bounty projects (currently anonymous) on Code4rena. We invited audit experts to verify our findings. In the end, we found 6 critical vulnerabilities, which were recognized by the project team or audit experts. In particular, one vulnerability was not discovered by any tools, marked as a great finding. This demonstrates the real-world value of TrustLLM. Due to page limitations, we have included these case studies in the supplementary materials for interested readers.
@@ IV-C RQ1 - Performance Comparison
Firstly, we compared TrustLLM with LLMs based on zero-shot learning, as shown in Table I. Our method also uses a zero-shot approach during the inference phase. We considered two proprietary models (GPT-4 and GPT-3.5) and three open-source models (CodeLlama-13b and CodeLlama-34b). For the open-source models, we strictly adhered to their prompt formats. Huggingface Transformer [58] has integrated these prompt formats into its framework. The format conversion is completed by calling apply_chat_template. CodeLlama requires adding [INST] and [/INST] as well as special tags <<SYS>> and <</SYS>>. As shown in Table I, after fine-tuning, our proposed strategy significantly outperforms the baseline models in the zero-shot scenario in terms of F1, accuracy, and precision, achieving high scores of $0.9121$, $0.8934$, and $0.9111$, respectively. However, when examining the recall, we notice that, the baseline models all performed excellently. Notably, GPT-4 and GPT-3.5 achieved a recall score of $1$. We checked the confusing matrix and found that all test data are labelled by the vulnerability. For GPT-4 and GPT-3.5, we adopted the prompts which are provided by our industrial partner, a Web3 security company.
TABLE I: Performance comparison between TrustLLM’s Detector and zero-shot LLMs.
@@
TrustLLM 0.9121 0.8934 0.9316 0.9111
TABLE II: Performance comparison between TrustLLM’s Detector and other fine-tuned models.
@@
As shown in Table II, TrustLLM achieves the highest scores of F1, Recall, and Accuracy among all methods, $0.9121$, $0.8934$ and $0.9111$. In terms of precision, CodeT5 achieves the best score $0.9431$. CodeLlama-13b-class is second only to our method regarding vulnerability detection rate, and the performance is relatively close. GraphCodeBERT and UnixCoder perform worse than CodeLlama-13b-class. Although CodeT5 achieves the highest precision, its other metrics are lower than GraphCodeBERT and UnixCoder. CodeBERT has the worst performance. Additionally, the accuracy scores of these models are relatively high (all are more than $0.91$), indicating that many of the predicted risky vulnerabilities are indeed risky.
Answer for RQ1: The performance of TrustLLM’s Detector exceeds that of traditional full-model fine-tuning, LoRA fine-tuning in a classification manner, and LLMs based on in-context learning. The performance of fine-tuned models is also better than that of zero-shot learning.
@@ IV-D RQ2 - Explanation Alignment
Figure 7: Comparing TrustLLM with the integration models that make decisions and explain the vulnerabilities simultaneously.
@@
Answer for RQ3: TrustLLM achieved better detection performance than the integration model that outputs labels and reasons simultaneously. We confirmed that the model struggles to focus on the labels when required to output both types of information, as evidenced by our inclusion of label-only data in the verification process.
@@ IV-F RQ4 - Effectiveness of Majority Voting
Our research explored a method using multiple prompts and a voting mechanism for Detector to determine the final label. This method aims to enhance the model’s precision and credibility. During the evaluation process, we continued to use metrics such as the F1 score, recall, precision, and accuracy. We calculated these metrics for each prompt individually for comparative analysis, as shown in Table III. It should be noted that the first row Single-prompt indicates that we used only one prompt format to train Detector. Prompt-1, Prompt-2, Prompt-3, Prompt-4, and Prompt-5 represent the results for each prompt after multiple-prompt training. The last row shows the results after majority voting, indicating that majority voting can improve the overall performance of TrustLLM, with both the F1 score and accuracy being the highest. At the same time, except for Single-prompt, we noticed minimal performance differences among multiple prompts. Single-prompt performed much worse than the others. Training with multiple prompts can improve model performance compared to using only one prompt during training.
@@ IV-G RQ5 - Impact of Additional Information
Through comparative experiments as shown in Table IV, we found that this calling contextual information did not improve the model’s overall performance. In the second row, Call, we used the prompts with the calling information and then employed majority voting to decide the prediction result. For the third row, Call-OutCall, we used prompts both with and without calling information and also used majority voting. Compared with TrustLLM’s Detector, they exhibited lower precision, accuracy, and f1, with almost the same recall.
@@ V Related Work
Vulnerability detection has been critical problems for the healthy and sustainable development of the software ecosystem, especially on blockchain and smart contracts. Traditional vulnerability detection methods, such as those based on predefined static analysis rules [3], often lack good generalization capabilities and are difficult to extend to new types of vulnerabilities. In addition, some logic-related vulnerabilities [2] are also difficult to abstract into static analysis rules. Researchers have employed a deep learning-based approach to address this issue. For example, Zhuang et al. [59] used graph neural networks to detect vulnerabilities in smart contracts. Liu et al. [60] fused the interpretable graph features with expert patterns to achieve better results and interpretable weights. Wu et al. [61] utilized pre-training technique and crucial data flow graphs for the detection of smart contract vulnerabilities.
With the advent of large language models (LLMs), researchers are utilizing not only traditional deep learning models but also LLMs for vulnerability detection. For example, Ullah et al. [62] evaluated LLMs on vulnerability detection tasks and found that they may not perform well. Fu et al. [63] further analyzed the gap for LLMs in detecting vulnerabilities. Thapa et al. [64] leveraged LLMs for software vulnerability detection, and David et al. [12] used LLMs for smart contract vulnerability tasks. Alqarni et al. [65] fine-tuned the BERT model [66] for source code vulnerability detection. Sun et al. [15] conducted an empirical study to find out how to enhance the ability of LLMs to detect vulnerabilities. In addition, Mathews et al. [67], Hu et al. [14], and Purba et al. [68] made efforts to utilize LLMs for detecting vulnerabilities. Some research also fused large language models with tradition program analysis methods. Sun et al. [13] proposed GPTScan for smart contracts, leveraging static program analysis to reduce the false positives of LLMs. Li et al. [69] proposed LLift for integrating LLMs with static analysis tools.
@@ VI Threats to Validity
In the data collection process, there is a risk of data bias, which might prevent models trained and tested on these data from generalizing accurately. Moreover, the precision of data labelling significantly impacts model performance. To mitigate these issues, we collected verified data from real and public audit reports and utilized the latest tools, such as GPTScan [13] and LLM4Vuln [15], to assist in cleaning and annotating the data. It is important to note that external links in the data could induce LLM to produce incorrect information; therefore, we performed data cleaning to remove these links. Considering LLMs’ sensitivity to input data, we standardized the code data, including removing unnecessary spaces without changing code semantics, to enhance the model robustness and reliability. To maximize the performance of the zero-shot learning of GPT-3.5 and GPT-4, we adopted and optimized the prompts from our partner, a Web3 security company. These prompts have been integrated into their working pipeline. For open-source models, we collaborated with an auditing expert to adapt their prompts for these models.
@@
VII Conclusion
In this paper, we proposed TrustLLM, the first smart contract auditing framework that combines fine-tuning and LLM-based agents to detect vulnerabilities and explain the results. We adopted a multiple-prompt-based strategy and applied LoRA-based fine-tuning to train the Detector and Reasoner. The former generates results based on a majority voting mechanism, while the latter provides multiple alternative explanations based on different inference paths. Furthermore, we introduced two LLM agents, Ranker and Critic, to collaborate in selecting the most appropriate explanation. Our approach demonstrated superior performance in zero-shot scenarios compared to zero-shot LLM learning and traditional full-model fine-tuning methods. We studied the performance improvement brought by the majority voting strategy and compared different LoRA training methods, providing the rationality of our choice. We also explored how additional calling context affects our model’s performance. For future work, we will focus on enhancing the model’s stability and its alignment with human preferences.

### v2

@@ IV-A Experimental Setup
In our study, we carefully selected a series of benchmark models, categorized into two groups: LLMs for zero-shot learning and pre-trained code models based on fine-tuning, to ensure a comprehensive and sound comparative analysis. For zero-shot learning LLMs, we chose CodeLlama-13b-Instruct, CodeLlama-34b-Instruct [49], GPT-3.5, and GPT-4 as benchmarks, representing the current state-of-the-art. Additionally, we selected CodeBERT [23], GraphCodeBERT [58], CodeT5 [28], UnixCoder [59], and CodeLlama-13b [49] to train classifiers. Among these, CodeBERT, GraphCodeBERT, CodeT5, and UnixCoder underwent a complete model fine-tuning process to adapt to the specific code classification task. In particular, CodeLlama-13b employs LoRA for lightweight tuning and uses the last token representation for classification. Note that that our method is different; iAudit’s Detector achieves classification by generating label names as task outputs.
@@ IV-B Research Questions (RQs)
How does the performance of iAudit in detecting vulnerabilities compare to other models? This question aims to understand how the effectiveness of Detector in detecting vulnerabilities compares to that of other existing models. The focus is on comparative analysis, involving metrics accuracy, precision, recall, and F1 score, to evaluate and contrast the performance.
@@
To what extent do the explanations generated by iAudit’s Reasoner align with the real reasons? RQ2 concerns the quality of the explanations the Reasoner provided for the decision of the Detector. It questions whether the reasons given by iAudit correspond to the actual reasons behind the vulnerabilities, emphasizing the interpretability and trustworthiness of the model.
@@
How does iAudit compare with an integration model that performs detection and reasoning simultaneously? Our method is based on a generative model, with two models trained on the generated labels and reasons, respectively. Another approach uses a single model to generate both reasons and labels. This question explores the effectiveness and impact of integrating the Detector and Reasoner components into one.
@@
Besides the RQs above, we also used our model to audit two bounty projects (currently anonymous) on Code4rena. We invited audit experts to verify our findings. In the end, we found 6 critical vulnerabilities, which were recognized by the project team or audit experts. In particular, one vulnerability was not discovered by any tools, marked as a great finding. This demonstrates the real-world value of iAudit. Due to page limitations, we have included these case studies in the supplementary materials for interested readers.
@@ IV-C RQ1 - Performance Comparison
Firstly, we compared iAudit with LLMs based on zero-shot learning, as shown in Table I. Our method also uses a zero-shot approach during the inference phase. We considered two proprietary models (GPT-4 and GPT-3.5) and three open-source models (CodeLlama-13b and CodeLlama-34b). For the open-source models, we strictly adhered to their prompt formats. Huggingface Transformer [60] has integrated these prompt formats into its framework. The format conversion is completed by calling apply_chat_template. CodeLlama requires adding [INST] and [/INST] as well as special tags <<SYS>> and <</SYS>>. As shown in Table I, after fine-tuning, our proposed strategy significantly outperforms the baseline models in the zero-shot scenario in terms of F1, accuracy, and precision, achieving high scores of $0.9121$, $0.8934$, and $0.9111$, respectively. However, when examining the recall, we notice that, the baseline models all performed excellently. Notably, GPT-4 and GPT-3.5 achieved a recall score of $1$. We checked the confusing matrix and found that all test data are labelled by the vulnerability. For GPT-4 and GPT-3.5, we adopted the prompts which are provided by our industrial partner, MetaTrust Labs, a Web3 security company.
TABLE I: Performance comparison between iAudit’s Detector and zero-shot LLMs.
@@
iAudit 0.9121 0.8934 0.9316 0.9111
TABLE II: Performance comparison between iAudit’s Detector and other fine-tuned models.
@@
As shown in Table II, iAudit achieves the highest scores of F1, Recall, and Accuracy among all methods, $0.9121$, $0.8934$ and $0.9111$. CodeLlama-13b-class is second only to our method regarding vulnerability detection rate, and the performance is relatively close. GraphCodeBERT and UnixCoder perform worse than CodeLlama-13b-class. Although CodeT5 achieves the highest precision at $0.9431$, its other metrics are lower than GraphCodeBERT and UnixCoder. CodeBERT has the worst performance. Additionally, the accuracy scores of these models are relatively high (all are more than $0.91$), indicating that many of the predicted risky vulnerabilities are indeed risky.
Regarding precision, iAudit produce more false positives than CodeT5, GraphCodeBERT, and CodeBERT. This is primarily because iAudit uses a longer context length (16k) compared to these models, whose 512-length context truncation causes the vulnerable and safe samples to align more closely in length, thus affecting the apparent code complexity. However, our dataset shows that vulnerable code is relatively more complex than safe code. This leads iAudit to output 10 unique false positives with more complex code.
Answer for RQ1: The performance of iAudit’s Detector exceeds that of traditional full-model fine-tuning, LoRA fine-tuning in a classification manner, and LLMs based on in-context learning. The performance of fine-tuned models is also better than that of zero-shot learning.
@@ IV-D RQ2 - Explanation Alignment
Figure 7: Comparing iAudit with the integration models that make decisions and explain the vulnerabilities simultaneously.
@@
Answer for RQ3: iAudit achieved better detection performance than the integration model that outputs labels and reasons simultaneously. We confirmed that the model struggles to focus on the labels when required to output both types of information, as evidenced by our inclusion of label-only data in the verification process.
@@ IV-F RQ4 - Effectiveness of Majority Voting
Our research explored a method using multiple prompts and a voting mechanism for Detector to determine the final label. This method aims to enhance the model’s precision and credibility. During the evaluation process, we continued to use metrics such as the F1 score, recall, precision, and accuracy. We calculated these metrics for each prompt individually for comparative analysis, as shown in Table III. It should be noted that the first row Single-prompt indicates that we used only one prompt format to train Detector. Prompt-1, Prompt-2, Prompt-3, Prompt-4, and Prompt-5 represent the results for each prompt after multiple-prompt training. The last row shows the results after majority voting, indicating that majority voting can improve the overall performance of iAudit, with both the F1 score and accuracy being the highest. At the same time, except for Single-prompt, we noticed minimal performance differences among multiple prompts. Single-prompt performed much worse than the others. Training with multiple prompts can improve model performance compared to using only one prompt during training.
@@ IV-G RQ5 - Impact of Additional Information
Through comparative experiments as shown in Table IV, we found that this calling contextual information did not improve the model’s overall performance. In the second row, Call, we used the prompts with the calling information and then employed majority voting to decide the prediction result. For the third row, Call-OutCall, we used prompts both with and without calling information and also used majority voting. Compared with iAudit’s Detector, they exhibited lower precision, accuracy, and f1, with almost the same recall.
@@ V Related Work
Vulnerability detection has been a critical issue for the healthy and sustainable development of the software ecosystem, especially in blockchain and smart contracts. Traditional vulnerability detection methods, such as those based on predefined static analysis rules [3], often lack robust generalization capabilities and are difficult to adapt to new types of vulnerabilities. Moreover, some logic-related vulnerabilities [2] are also challenging to encapsulate into static analysis rules. To address this issue, researchers have employed deep learning-based approaches. For example, Zhuang et al.[61] used graph neural networks to detect vulnerabilities in smart contracts. Liu et al.[62] combined interpretable graph features with expert patterns to achieve better results and interpretable weights. Wu et al. [63] utilized a pre-training technique and critical data flow graphs for the detection of smart contract vulnerabilities.
With the advent of large language models, researchers are not only utilizing traditional deep learning models but also LLMs for vulnerability detection. For example, Ullah et al. [64] evaluated LLMs on vulnerability detection tasks and found that they may not perform well. Fu et al. [65] further analyzed the gap for LLMs in detecting vulnerabilities. Thapa et al. [66] leveraged LLMs for software vulnerability detection, and David et al. [12] used LLMs for smart contract vulnerability tasks. Alqarni et al. [67] fine-tuned the BERT model [68] for source code vulnerability detection. Sun et al. [15] proposed an unified evaluation framework, LLM4Vuln, to enhance the ability of LLMs to detect vulnerabilities. Some research also fused large language models with traditional program analysis methods. Sun et al. [13] proposed GPTScan for smart contracts, leveraging static program analysis to reduce the false positives of LLMs. Li et al. [69] proposed LLift for integrating LLMs with static analysis tools. SmartInv [70] and PropertyGPT [71] further integrated large language models with formal verification methods, aiming not only to detect vulnerabilities but also to prove that a piece of code is secure.
@@ VI Threats to Validity
In the data collection process, there is a risk of data bias, which might prevent models trained and tested on these data from generalizing accurately. Moreover, the precision of data labelling significantly impacts model performance. To mitigate these issues, we collected verified data from real and public audit reports and utilized the latest tools, such as GPTScan [13] and LLM4Vuln [15], to assist in cleaning and annotating the data. It is important to note that external links in the data could induce LLM to produce incorrect information; therefore, we performed data cleaning to remove these links. Considering LLMs’ sensitivity to input data, we standardized the code data, including removing unnecessary spaces without changing code semantics, to enhance the model robustness and reliability. To maximize the performance of the zero-shot learning of GPT-3.5 and GPT-4, we adopted and optimized the prompts from our partner, MetaTrust Labs. These prompts have been integrated into their working pipeline. For open-source models, we collaborated with an auditing expert to adapt their prompts for these models.
@@
VII Limitation
Although our method performed excellently in trials, its primary advantage lies in detecting logic vulnerabilities in smart contracts, which account for more than 80% of exploitable vulnerabilities according to a recent study [2]. As such, while iAudit’s training data does include instances of Reentrancy, Overflow, and Underflow vulnerabilities, handling them is not the major usage scenario of iAudit. Indeed, these traditional contract vulnerabilities are theoretically more suitable for detection by program analysis methods. That said, both AI-based and PL-based methods have their unique comfort zones, and since iAudit is fully AI-based, this paper compared it with other AI-based methods only. Additionally, despite our efforts to mitigate the phenomenon of hallucinations in large language models (LLMs) through voting and agents, hallucinations may still occur. Finally, since our tools rely on LLMs, there are certain hardware requirements. While the models can be compressed to run on less powerful GPUs, this compression may result in reduced model performance.
VIII Conclusion
In this paper, we proposed iAudit, the first smart contract auditing framework that combines fine-tuning and LLM-based agents to detect vulnerabilities and explain the results. We adopted a multiple-prompt-based strategy and applied LoRA-based fine-tuning to train the Detector and Reasoner. The former generates results based on a majority voting mechanism, while the latter provides multiple alternative explanations based on different inference paths. Furthermore, we introduced two LLM agents, Ranker and Critic, to collaborate in selecting the most appropriate explanation. Our approach demonstrated superior performance in zero-shot scenarios compared to zero-shot LLM learning and traditional full-model fine-tuning methods. We studied the performance improvement brought by the majority voting strategy and compared different LoRA training methods, providing the rationality of our choice. We also explored how additional calling context affects our model’s performance. For future work, we will focus on enhancing the model’s stability and its alignment with human preferences.
Acknowledgment
We thank all the reviewers for their detailed and constructive comments. We also express our gratitude to all colleagues at MetaTrust Labs for their assistance in deploying the TrustLLM’s iAudit model. This research/project is supported by the National Research Foundation, Singapore, and the Cyber Security Agency under its National Cybersecurity R&D Programme (NCRP25-P04-TAICeN), the National Research Foundation, Singapore, and DSO National Laboratories under the AI Singapore Programme (AISG Award No: AISG2-GC-2023-008), and NRF Investigatorship NRF-NRFI06-2020-0001. Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of National Research Foundation, Singapore and Cyber Security Agency of Singapore. Daoyuan Wu was also partially supported by an HKUST grant.